OpenStudy (anonymous):

write the difference of 48-(-19) as a sum than simplify?

OpenStudy (anonymous):

62???

OpenStudy (anonymous):

48-(-19)=48+19= 67

OpenStudy (anonymous):

i know its 67 but its not in the choices

OpenStudy (anonymous):

but if you simplify its 62 right?

OpenStudy (anonymous):

What is your full question?

OpenStudy (anonymous):

no wonder why i got it wrong cuz i thought the -14 was a -19 sryy

OpenStudy (anonymous):

no prob